OpenStudy (mattduhhelper):

The price of a pair of shoes is marked down 33%. The original price was $59.85. What is the sales price of the shoes? $16.85 $19.75 $40.10 $45.00

OpenStudy (mattduhhelper):

@Awolflover1

Awolflover1 (awolflover1):

59.85-33%=? @563blackghost am i correct? :^)

563blackghost (563blackghost):

Im so sorry @Awolflover1 I was off. Since the shoes are `33%` off we would multiply the origional price by the percentage. `p = how much money taken from original price` \(\huge\bf{59.85 \times .33 =p}\) Once you find that you would subtract by original price. \(\huge\bf{59.85 - p = \color{red}{answer}}\)

Awolflover1 (awolflover1):

Its okay :3 <3